OpenStudy (anonymous):

Determine the solution of the system. This vertically set up.Help!!!. y=x+6 y=2x

OpenStudy (anonymous):

y=x+6 2x=x+6 x=6 y=12

OpenStudy (anonymous):

thanks so much :)